OpenStudy (anonymous):

Why can this solid of revolution problem not be solved with shell? (typing question w/ eqation maker)

OpenStudy (anonymous):

\[For the formula:2\pi \int\limits\limits_{a}^{b}phdx\] Where a = 0, b=4 p=x and h=\[\sqrt{x}\]

OpenStudy (anonymous):

After plugging in and evaulating I am getting \[128\pi/5\], my texbook answer is \[32\pi/5\] Am i making an error or is shell method just impossible in this scenario? If it is not possible, please explain why, thanks!

OpenStudy (anonymous):

Please write the question as it was asked. That way we take a fresh take, a fresh look at it.

OpenStudy (anonymous):

Using the shell method, find the volume of: \[\sqrt(x)\] with bounds [0,4] after it has been revolved around the y axis.

OpenStudy (anonymous):

and the bounds on the y is 0,2, but that is not relevant to the formula

OpenStudy (anonymous):

(y-axis that is)

OpenStudy (anonymous):

It is confusing '...and the bounds of y is 0,2' what is that info.

OpenStudy (anonymous):

It just means there is a line at y=2

OpenStudy (anonymous):

Is that in the question, or is that something you figured out?

OpenStudy (anonymous):

if there wasn't it wouldnt make a shape when revolved cause the top would be endless

OpenStudy (anonymous):

There is actually a picture, hold on one second

OpenStudy (anonymous):

If you can imagine these graphs zoomed out more: http://www.wolframalpha.com/input/?i=y%3Dsqrt%28x%29%2C+y%3D2

OpenStudy (anonymous):

But what is the actual question? You wrote the question and then you added something. It is that part of the question.

OpenStudy (anonymous):

The actual question is exactly what I wrote only it asked to solve with disk method

OpenStudy (anonymous):

Using the shell method, find the volume of: (√x) with bounds [0,4] after it has been revolved around the y axis.

OpenStudy (anonymous):

only using disk method

OpenStudy (anonymous):

Disk or shell?

OpenStudy (anonymous):

It askd for disk, I am trying shell

OpenStudy (anonymous):

But I got an incorrect answer so I am tring to figure out if shell does not work or if i messed up

OpenStudy (anonymous):

Ok, here is an illustration: http://www.scriblink.com/index.jsp?act=phome&roomid=1650&KEY=9445365509101FCAF650C4612FFA0010

OpenStudy (anonymous):

Any way, your answer is not far off. Just write your work on your whiteboard and we will see if you made a little arithmetic or algebra error.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Done

OpenStudy (anonymous):

You checking?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I know that shell method is not the correct method in some cases but I do not know why

OpenStudy (anonymous):

So I am trying to see if this is one of those cases so I can figure out why

OpenStudy (anonymous):

I get 128pi/5, but if I understand you it asked for disk method, did you get it right in disk method?

OpenStudy (anonymous):

yea, I got it right with disk

OpenStudy (anonymous):

32pi/5

OpenStudy (anonymous):

Do we have a pickle?

OpenStudy (anonymous):

I'm checking.

OpenStudy (anonymous):

I think you have to convert it in terms of y. And then use some y. In your diagram you used some x. It is good that you investigate different techniques you would be very successful. When you have time change f(x) to g(y) and so on.

OpenStudy (anonymous):

Ah yea, I think that may be it

OpenStudy (anonymous):

I'll give that a try, thanks

OpenStudy (anonymous):

It wont let me give you a medal?